OpenStudy (anonymous):

School books are printed on a four color web offset printing press a cylinder on a press has a 13.37 inch diameter. The linear speed of a point on the cylinder's surface is 16.53 feet per second. What is the angular speed of the cylinder in revolutions per hour? Printers often refer to the angular speed as impressions per hour or IPH. HELP!

OpenStudy (anonymous):

@genius12

OpenStudy (davidusa):

hi what art of math is this?

OpenStudy (anonymous):

angular/ linear velocity

OpenStudy (davidusa):

which is... physics or what

OpenStudy (anonymous):

umm well if you dont know how to do it. its ok just say so

OpenStudy (anonymous):

its math

OpenStudy (davidusa):

lets see. let me look longer and carefully

OpenStudy (davidusa):

idk

OpenStudy (davidusa):

draw a diagram- it makes it easier to understand the question. i will summon a bunch of my friends.

OpenStudy (anonymous):

ok thnks for trying i guess

OpenStudy (davidusa):

@Homeworksucks @TheForbiddenFollower @wolfe8 @ducati289 @habibmatatta

OpenStudy (davidusa):

just random people except for wolfe8- i know her

OpenStudy (davidusa):

@ducati289 @ybarrap @aurorablue01 @Lethal i know aurorablue

OpenStudy (anonymous):

Oh Em Gee

OpenStudy (davidusa):

hi wolfe thanks for coming thanks letal

OpenStudy (anonymous):

We need to do some DA here

OpenStudy (davidusa):

what?

OpenStudy (anonymous):

Dimensional Analysis

OpenStudy (davidusa):

oh ok

OpenStudy (anonymous):

yup thats it

OpenStudy (wolfe8):

The formula for angular speed is \[\omega =\frac{ v }{ r }\]

OpenStudy (davidusa):

@esshotwired @RadEn @WinterRaine @surjithayer

OpenStudy (anonymous):

@DavidUsa

OpenStudy (anonymous):

@wolfe8 no clue what youre talkin abt........

OpenStudy (davidusa):

what do you want @Lethal @Lethal @Lethal @Lethal @Lethal @Lethal @Lethal @Lethal @Lethal

OpenStudy (anonymous):

That's the formula for Angular Velocity which is what we're trying to find in rph

OpenStudy (anonymous):

Right now we're in seconds so we also have to do some DA

OpenStudy (anonymous):

can u like draw it so i can actually see what the ratios look like?

OpenStudy (wolfe8):

Try reading this first: http://hyperphysics.phy-astr.gsu.edu/hbase/rotq.html

OpenStudy (davidusa):

i dont know anything about this stuff so please dont mention me with the @ symbol

OpenStudy (anonymous):

this is what ive got so far. tell me if its right??

OpenStudy (anonymous):

Dude perl already helped you out with this he said: w = (2 * 16.53 ) / 13.37 * rad/sec * rev/ (2pi * rad) * 3600 sec/ 1hr

OpenStudy (davidusa):

bump this question up music101

OpenStudy (anonymous):

This is very good for your matter http://people.wku.edu/david.neal/117/Unit2/AngVel.pdf

OpenStudy (anonymous):

already read that

OpenStudy (anonymous):

im just asking if what i have up there is all i need

OpenStudy (anonymous):

c=pi d=pi*13.37/12=3.5 feet time taken to complete one revolution=3.5/16.53 =0.21175 no. of revolutions per hour=/0.21175*3600=17001 revolutions per hour

OpenStudy (anonymous):

ok thanks